A circle with radius 10x is inscribed in a square.
part a: Write an expression for the area of the circle.
part b: Write an expression for the area of the square.
part c: Form a simplified rational expression from the ratio of the two areas. Use the circle area as the numerator.

Answer by stanbon(75887)

A circle with radius 10x is inscribed in a square.
Sketch that picture; label the radius as 10x; then the diagonal = 20x
Then each side is 20x/sqrt(2) = 10sqrt(2)x
---------
part a: Write an expression for the area of the circle.
Area = pi*radius^2 = pi*100x^2
-----------------------------------------
part b: Write an expression for the area of the square.
Area = side^2 = 100*2x^2 = 200x^2
-------
part c: Form a simplified rational expression from the ratio of the two areas. Use the circle area as the numerator.
Ans: [pi*100x^2] / [200x^2] = pi/2
------------
Cheers,
Stan H.
------------